OpenStudy (anonymous):

simplify the following fractions completely: (x(x+1))/(x^2)

OpenStudy (anonymous):

help pls

OpenStudy (agreene):

\[\frac{x+1}{x}\]

jhonyy9 (jhonyy9):

x(x+1) x+1 ------ = ----- x*x x

OpenStudy (anonymous):

2a div a^2b

jhonyy9 (jhonyy9):

will be 2/ab

OpenStudy (anonymous):

\[9x + 3 \div 3x\]

jhonyy9 (jhonyy9):

us parantheses please

OpenStudy (anonymous):

(9x\[(9x+3) \div (3x)\]

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

help asap

jhonyy9 (jhonyy9):

9x/3x +3/3x = 3+1/x

jhonyy9 (jhonyy9):

or (3x+1)/x

OpenStudy (anonymous):

\[(4a + 5a ^{2}) \div (5a)\]

OpenStudy (anonymous):

help pls asap

jhonyy9 (jhonyy9):

(4/5) +a

jhonyy9 (jhonyy9):

so give me medals for

OpenStudy (anonymous):

\[((2x)^{2} - (8x)) \div (4x)\]

jhonyy9 (jhonyy9):

x-2